# Recommendation and Temptation

### v1

@@ Abstract.
Traditional recommender systems based on utility maximization and revealed preferences often fail to capture users’ dual-self nature, where consumption choices are driven by both long-term benefits (enrichment) and desire for instant gratification (temptation). Consequently, these systems may generate recommendations that fail to provide long-lasting satisfaction to users. To address this issue, we propose a novel user model that accounts for this dual-self behavior and develop an optimal recommendation strategy to maximize enrichment from consumption. We highlight the limitations of historical consumption data in implementing this strategy and present an estimation framework that makes minimal assumptions and leverages explicit user feedback and implicit choice data to overcome these constraints. We evaluate our approach through both synthetic simulations and simulations based on real-world data from the MovieLens dataset. Results demonstrate that our proposed recommender can deliver superior enrichment compared to several competitive baseline algorithms that assume a single utility type and rely solely on revealed preferences. Our work emphasizes the critical importance of optimizing for enrichment in recommender systems, particularly in temptation-laden consumption contexts. Our findings have significant implications for content platforms, user experience design, and the development of responsible AI systems, paving the way for more nuanced and user-centric recommendation approaches.
@@ 1. Introduction
Consumption of online content has become a central fixture in everyday life, spanning various forms such as articles, videos, films, social media, and more. This consumption accounts for a substantial portion of people’s time, significantly impacting daily routines and leisure activities. For example, according to Statista [^1] https://www.statista.com/statistics/1359403/us-time-spent-per-day-netflix-tiktok-youtube/, adults in the United States spend an average of 46 minutes on YouTube, 48 minutes on TikTok, and 61 minutes on Netflix every day. These statistics underscore the pervasive nature of digital content in modern life and highlight the importance of efficient content discovery mechanisms. While recommender systems help users navigate this vast sea of content more effectively, there is growing concern that these systems may be hindering users from getting the maximum benefit possible out of consumption. This potential shortcoming raises questions about the fundamental assumptions underlying recommender system design and their impact on user experience.
These limitations stem from two core premises of the current recommender design paradigm:
@@
    Each user has a utility function that represents the benefit they derive from a given option.
@@
These assumptions, while providing a foundation for many recommender systems, may oversimplify the complex nature of human decision-making in content consumption. As a result, they could lead to recommendations that fail to account for the nuanced ways in which users interact with and benefit from digital content.
In this paper, we present a novel approach to recommender design by challenging the second premise: we posit that while users possess a utility function representing derived value or long-term benefit, their decision-making process is influenced by additional content dimensions, resulting in consumption choices that seemingly contradict their utility-maximizing nature. This departure from the traditional utility maximization principle finds strong support in existing literature. For instance, Daniel Kahneman (Kahneman, 2011) distinguishes between two types of happiness: experienced happiness, which encompasses immediate emotions, and life satisfaction or evaluative happiness, representing a long-term perspective on well-being. Similarly, several studies (Thaler and Shefrin, 1981; Kleinberg et al., 2022) conceptualize users as a combination of short-term and long-term selves, with the former prioritizing immediate gratification and the latter optimizing for long-term benefits from consumption. Drawing inspiration from these conceptual frameworks, we identify two distinct dimensions of content that influence consumption choices:
@@
    Temptation quantifies the instant gratification provided by content, analogous to Kahneman’s experienced happiness or short-term benefits in dual-self models.
@@
    Enrichment measures the long-term value or benefit offered by content, similar to Kahneman’s life satisfaction or long-term benefits in dual-self models.
Empirical observations of online user behavior substantiate the distinction between these concepts. A notable example is the phenomenon observed on platforms like YouTube, where users feel “sucked down a rabbit hole” and engage in prolonged consumption of similar content they themselves acknowledge as mediocre (Woolley and Kenyon, 2023). This behavior exemplifies the tension between temptation and enrichment: users find themselves drawn to and unable to disengage from highly tempting content, despite recognizing its limited long-term value or enrichment potential.
The distinction between these concepts identifies a critical blind spot in traditional recommender systems. When temptation influences consumption choices, historical data ceases to be a pure reflection of enrichment. Consequently, recommenders relying solely on this data, without an appropriate user model, struggle to differentiate between enrichment and temptation. Furthermore, recommenders that fail to account for temptation and its impact on consumption choices in their objective cannot ensure high enrichment from consumption. For instance, a recommender driven exclusively by engagement metrics may reinforce temptation, inadvertently preventing users from selecting high-enrichment content both on-platform and off-platform. Conversely, a recommender that simply ranks items by enrichment, without considering actual user decision-making processes, may fail to engage users, potentially driving them towards off-platform options with high temptation but low enrichment. This dichotomy underscores the critical importance of integrating both enrichment and temptation considerations in the design of effective recommender systems.
1.1. Contributions
This paper makes several key contributions to the field of recommender systems:
@@ 1.1. Contributions
    First, we introduce a novel consumption model that captures user decision-making by considering both enrichment and temptation, providing a more comprehensive understanding of user behavior.
@@
    Second, we formulate an appropriate objective function for recommender systems aimed at maximizing user enrichment, and describe the optimal recommendation strategy to achieve this goal.
@@
    Third, we propose and implement reasonable assumptions to address the limitations of historical data, and develop a basic framework for estimating crucial user and item information, demonstrating the feasibility of a practical implementation of our proposed strategy.
The remainder of this paper is structured as follows: Section 2 reviews existing literature on consumption under temptation and recommender systems. Section 3 details our proposed consumption model. Section 3.1 explains our recommendation objective and optimal strategy. Section 4 outlines our estimation framework and necessary assumptions. Section 5 presents an agent-based simulation comparing our approach to baseline recommendations. Finally, Section 6 evaluates our algorithm using the MovieLens dataset, comparing it against ratings-based and behavior-based algorithms.
2. Related Work
Our work synthesizes insights from behavioral studies of users under temptation and recommender system design.
User Behavior Under Temptation.
The prevalence of temptation and the human tendency to prioritize instant gratification over long-term benefits is well-documented across various domains. Experimental studies have examined this phenomenon in contexts ranging from purchase behaviors (Wertenbroch, 1998) and appointment compliance (Trope and Fishbach, 2000) to smoking cessation (Giné et al., 2010), savings (Thaler and Benartzi, 2004), and study concentration (Ariely and Wertenbroch, 2002). In the digital content sphere, Milkman et al. (Milkman et al., 2009) provided evidence of temptation-driven behavior in film consumption through a field experiment with QuickFlix users.
Economists have proposed several theoretical models to explain this behavior. One class of models, including those by Strotz (Strotz, 1955), Akerlof (Akerlof, 1991), Laibson (Laibson, 1997), and O’Donoghue and Rabin (O’Donoghue and Rabin, 1999), suggests that people demonstrate time-inconsistent preferences, prioritizing current benefits in a way that violates their past relative preferences. Another class of models, proposed by Thaler and Sherfin (Thaler and Shefrin, 1981), Kleinberg et al. (Kleinberg et al., 2022), and others, introduces the concept of dual selves: a forward-looking self that cares about long-term benefits and a myopic self that focuses on instant gratification, with the latter having greater control over decision-making. Finally, a third class of models, proposed by Gul and Pesendorfer (Gul and Pesendorfer, 2001, 2004), explains user behavior by introducing temptation as an additional aspect of options beyond intrinsic value or enrichment, demonstrating how it affects people’s commitment and decision making.
Despite their differences, these models share a fundamental insight: user consumption choices are influenced by the interplay between enrichment and temptation. One could argue that time-inconsistent preferences arise from the interaction between two selves, with the myopic self prioritizing temptation and exerting greater control over choices. Furthermore, the forward-looking self may optimize for enrichment, while the myopic self maximizes the combination of enrichment and temptation. We draw inspiration from this shared perspective in designing our model of user decision-making, which aims to capture comprehensive picture of content consumption in digital platforms.
Recommender System Design.
The traditional recommendation design paradigm is rooted in behaviorism, relying on observed user behavior to predict and guide future interactions. Initially implemented using matrix factorization techniques popularized by the 2006 Netflix challenge, modern recommenders now employ sophisticated machine learning approaches, including neural networks for user-item feature matching (He et al., 2017) and graph neural networks for capturing user-item interaction networks (Wang et al., 2019; He et al., 2020).
However, the adequacy of behaviorism in recommender design has been increasingly questioned. Ekstrand and Willemsen (Ekstrand and Willemsen, 2016) argue for active user participation to discern whether recommenders fulfill long- or short-term user desires. Kleinberg et al. (Kleinberg et al., 2022) emphasize the need to disentangle decisions made by short and long-term selves, cautioning against an overemphasis on instant gratification. Milli et al. (Milli et al., 2023a) provide evidence that engagement-focused recommenders may fall short in user satisfaction.
Recent research has focused on separating content value from engagement metrics. Milli et al. (Milli et al., 2021, 2023b) propose a latent variable model to detect item enrichment from implicit signals, while Chang et al. (Chang et al., 2023) suggest a Bayesian network approach to incorporate current user intent in behavior prediction. Our work extends these efforts by not only disentangling enrichment from engagement but also identifying temptation as a key driver of engagement. We distinguish between recommending high-enrichment items and ensuring high enrichment through recommendations, providing an optimal strategy for the latter.
3. User Model and Recommendation Design
@@
We model user behavior within the confines of a single platform, considering $m$ users and $n$ items, where $m\gg n$. As previously discussed, for each user $j=1,\ldots,m$, each item $i=1,\ldots,n$ has two attributes: i) Enrichment $u_{j}(i)$ represents the long-term value or benefit the item provides to the user, and ii) Temptation $v_{j}(i)$ measures the instant gratification or short-term appeal of the item to the user.
@@
where $f_{rating}$ is monotone and non-decreasing.
An important assumption in our model is that $\lambda^{F}_{j}\geq\lambda^{C}_{j}\hskip 5.69054pt\forall j=1,\ldots,m$. This means that users tend to place more weight on enrichment when providing feedback than when choosing what to consume. This assumption is grounded in the observation that rating an item typically involves more deliberate reflection from a user compared to the often more impulsive act of choosing what to consume. This aligns with psychological theories suggesting that people are more likely to consider long-term value when engaged in reflective thought processes.
3.1. Recommender Design
@@
We then induct on the number of rounds (the depth of the tree), $T$. The intuition for the proof here is that if a recommendation strategy $\mathcal{A}$ is optimal, it must be locally greedy after the first round because of the induction hypothesis. Then, we can either swap the item recommended by $\mathcal{A}$ in round 1 with a later item without losing enrichment, or we can replace the item recommended by $\mathcal{A}$ in round 1 with an item that provides better enrichment in expectation. A detailed proof is provided in the appendix. ∎
@@ 4. Model Estimation
The optimal recommender strategy proposed in section 3.1 requires knowledge of both on-platform items and off-platform outside options for each user, specifically their enrichment and temptation values, as well as user-specific parameters. However, in reality, platforms do not possess this comprehensive information and must instead estimate these components from available historical data. To address this challenge, we first extend our model to elucidate the determination of enrichment and temptation values for items. Subsequently, we provide a framework for estimating these critical components and generating recommendations that align with our proposed strategy. This estimation process bridges the gap between theoretical optimality and practical implementation, enabling platforms to leverage our recommendation strategy effectively, even in the absence of perfect information. By doing so, we not only enhance the applicability of our model but also provide valuable insights into user behavior and content characteristics.

### v2

@@ Abstract.
Traditional recommender systems based on revealed preferences often fail to capture the fundamental duality in user behavior, where consumption choices are driven by both inherent value (enrichment) and instant appeal (temptation). Consequently, these systems may generate recommendations that prioritize short-term engagement over long-lasting user satisfaction. We propose a novel recommender design that explicitly models the tension between enrichment and temptation. We introduce a behavioral model that accounts for how both enrichment and temptation influence user choices, while incorporating the reality of off-platform alternatives. Building on this model, we formulate a novel recommendation objective aligned with maximizing consumed enrichment and prove the optimality of a locally greedy recommendation strategy. Finally, we present an estimation framework that leverages the distinction between explicit user feedback and implicit choice data while making minimal assumptions about off-platform options. Through comprehensive evaluation using both synthetic simulations and real-world data from the MovieLens dataset, we demonstrate that our approach consistently outperforms competitive baselines that ignore temptation dynamics either by assuming revealed preferences or recommending solely based on enrichment. Our work represents a paradigm shift toward more nuanced and user-centric recommender design, with significant implications for developing responsible AI systems that genuinely serve users’ long-term interests rather than merely maximizing engagement.
Recommender Systems, Online Platforms, Content Consumption, Behavioral Model, User Engagement, User Well-being
ccs: Information systems Recommender systemsjournalyear: 2025copyright: noneconference: Proceedings of the Nineteenth ACM Conference on Recommender Systems; September 22–26, 2025; Prague, Czech Republicbooktitle: Proceedings of the Nineteenth ACM Conference on Recommender Systems (RecSys ’25), September 22–26, 2025, Prague, Czech Republicdoi: 10.1145/3705328.3748063isbn: 979-8-4007-1364-4/2025/09
@@ 1. Introduction
Consumption of online content has become a central fixture in everyday life, spanning various forms such as articles, videos, films, social media, and more. This consumption accounts for a substantial portion of people’s time, significantly impacting daily routines and leisure activities. For example, according to Statista [^1] https://www.statista.com/statistics/1359403/us-time-spent-per-day-netflix-tiktok-youtube/, adults in the United States spend an average of 46 minutes on YouTube, 48 minutes on TikTok, and 61 minutes on Netflix every day. Recommender systems act as the guiding force during these excursions, helping users navigate the vast sea of content more effectively. The traditional recommendation design paradigm is rooted in behaviorism, relying on observed user behavior to predict and guide future interactions. Initially implemented using matrix factorization techniques popularized by the 2006 Netflix challenge, modern recommenders now employ sophisticated machine learning approaches, including neural networks for user-item feature matching (He et al., 2017) and graph neural networks for capturing user-item interaction networks (Wang et al., 2019; He et al., 2020) etc.
However, there is a growing concern that these systems may be guiding users to the wrong sets of content, therefore preventing them from getting the maximum benefit possible out of consumption. This potential shortcoming stems from two core premises of the current recommender design paradigm:
@@
    Each user has a single utility function that represents the benefit they derive from a given option.
@@
These assumptions, while providing a foundation for conventional recommenders, oversimplify the complex nature of human decision-making during content consumption.
1.1. Temptation
In this paper, we present an overhaul of recommender design by challenging these premises. Our first contribution is a novel behavioral model grounded in the literature on temptation and has the following three features:
@@ 1.1. Temptation
    Enrichment: Each item has some inherent value that the user derives from consuming it.
@@
    Temptation: Each item also has some inherent appeal that is orthogonal to its enrichment or inherent value, but nonetheless influences the consumption choices made by users. Note that the enrichment and temptation of an item is personalized for a given user, and can vary across users.
@@
    Outside Options: While the platform presents the user with a set of recommended items, the user has an additional outside option available to them. The user chooses their consumption from the recommended items and their available outside option according to a personalized weighted combination of enrichment and temptation.
The intuition is that while users would benefit from consuming highly enriching content, they may instead be tempted by options with inferior enrichment but high appeal, thus failing to act in their own best interest. This violates the first premise because users effectively operate with two utility functions: one representing their true preferences or end goals, and another representing their actual choice behavior. It also violates the second premise because users make consumption decisions by optimizing their choice behavior rather than their stated preferences, appearing to contradict the assumption of utility maximization.
The human tendency to do other than that which an individual wants is already observed by Paul in the Christian Bible, “I do not understand my own actions. For I do not do what I want, but I do the very thing I hate" (Romans 7:15). More recently, the research literature documents this mismatch across domains ranging from purchase behaviors (Wertenbroch, 1998) and appointment compliance (Trope and Fishbach, 2000) to smoking cessation (Giné et al., 2010), savings (Thaler and Benartzi, 2004), and study concentration (Ariely and Wertenbroch, 2002). In the context of digital media, Milkman et al. (Milkman et al., 2009) ran an empirical study on Quickflix and examined the orders in which DVDs were rented and returned, illustrating a difference between stated preferences (adding DVDs to the rental queue) and consumption (watch order of DVDs).
Economists have proposed several theoretical frameworks to explain this behavior. One class of models, including those by Strotz (Strotz, 1955), Akerlof (Akerlof, 1991), Laibson (Laibson, 1997), and O’Donoghue and Rabin (O’Donoghue and Rabin, 1999), suggests that people exhibit time-inconsistent preferences, prioritizing immediate benefits in ways that contradict their previously stated preferences. A second class of models, proposed by Thaler and Shefrin (Thaler and Shefrin, 1981) and others, introduces the concept of dual selves: a forward-looking self concerned with long-term benefits and a myopic self focused on instant gratification, with the latter exerting greater influence over actual decisions. A third class of models, developed by Gul and Pesendorfer (Gul and Pesendorfer, 2001, 2004), explains user behavior by treating temptation as an additional attribute of options beyond intrinsic value, showing how it influences commitment and decision-making processes.
Our model remains agnostic about the precise nature of enrichment and temptation. In our model, enrichment simply represents what users genuinely wish to optimize, whether their stated preferences, long-term benefits, or other intrinsic values. Temptation, in turn, captures users’ immediate desire to choose an item, independent of its enrichment value.
Our model implies that users’ observed behaviors from engagement may not serve as reliable indicators of their true preferences (e.g. revealed preferences) or derived value. Consequently, traditional recommenders that rely on behavioral signals and optimize for engagement metrics cannot effectively optimize for users’ stated desires or long-term well-being. Beginning with Ekstrand and Willemsen (Ekstrand and Willemsen, 2016), a growing body of literature has documented this fundamental limitation of engagement-based recommenders and often desiring instead to incorporate some form of explicit user feedback. Some scholars even characterize the continued reliance on revealed preferences as “indefensible” (Morewedge et al., 2023).
This aligns with qualitative evidence showing that users express hindsight regret when evaluating their experience on content platforms (McCrosky and Geurkink, 2021). Indeed, users on platforms like YouTube report feeling “sucked down a rabbit hole” and engaging in prolonged consumption of content they themselves acknowledge as mediocre (Woolley and Kenyon, 2023).
Milli et al. (Milli et al., 2023a) demonstrated in an empirical study on Twitter (now X) that users experienced reduced negative emotions and greater satisfaction from political tweets recommended based on stated preferences (e.g., all posts from those they follow) compared to actual engagement-based recommendations. Their results not only demonstrate the divergence between engagement-based recommendations and those based on stated preferences, but also reveal that engagement-based recommendations contain more anti-social content, highlighting potential societal harms beyond individual user enrichment loss.
Furthermore, recently Kleinberg et al. (Kleinberg et al., 2022) have used a dual-self model to illustrate how users become drawn to and are unable to disengage from highly tempting content, despite recognizing its limited long-term value or enrichment potential. Our behavioral model draws inspiration from theirs; however, ours is not based on a particular characterization of users’ inner selves. While both models are concerned with how tempting items may distort platform use, their model focuses on time-on-platform whereas ours focuses on which items the user chooses to consume. Additionally, we go beyond merely diagnosing user behavior to prescribe a recommender system that actively helps users overcome temptation.
1.2. Optimizing Enrichment
The literature discussed so far demonstrates the shortcoming of recommenders optimizing engagement metrics in helping users achieve their end goal, i.e. optimize their enrichment. An alternative is to simply try to reconstruct the enrichment of each item and recommend the items with the highest enrichment. However, in most realistic settings, this fails as illustrated in the following anecdote: one of the authors signed up for a Netflix subscription at the time when they shipped DVDs. The first DVD they ordered was “Lawrence of Arabia," a nearly four hour biopic on T. E. Lawrence. They returned it 4 months later, unwatched.
Outside options are also crucial to consider when optimizing enrichment. In realistic scenarios, content platforms are not insulated —they exist in an ecosystem with competing platforms and activities, framed as exogenous off-platform options available to a user when making consumption choices in our model. The author did finally watch “Lawrence of Arabia" on a 15-hour flight, where outside options were limited.
Industry leaders recognize the necessity of considering outside options in platform and recommender design. Luis von Ahn, creator of Duolingo, talks about “How to Make Learning as Addictive as Social Media” (von Ahn, 2023; Raz, 2020) as a necessity to combat the distraction of other apps on a user’s smartphone. He acknowledges navigating a tradeoff between the educational value of Duolingo and its ability to retain users. Similarly, Netflix CEO Reed Hastings talks about Netflix competing not just with other streamers but anything users do to “relax and unwind, hang out and connect” including sleep (Staff, 2017).
These challenges in optimizing enrichment arise from a fundamental insight captured by our model: recommended items need not resemble consumed items. There is a crucial distinction between recommended enrichment and consumed enrichment. While recommending highly engaging content is problematic, simply recommending enriching items may not be much better, as echoed by industry leaders who recognize this tension. Instead, platforms must carefully consider temptation when making recommendations. Moreover, platforms must exercise caution even when making multiple recommendations without outside options, as a single highly tempting item can deter users from selecting other recommendations regardless of their enrichment value.
Understanding users’ outside options is equally crucial for optimizing enrichment. Consider two contrasting scenarios: a user seeking quality content during leisure time that would otherwise be spent on low-value activities like doom scrolling, versus a student procrastinating on important exam preparation. In the first case, the platform should recommend sufficiently engaging content to retain the user while maximizing enrichment within this constraint. In the second case, where any platform content is likely less enriching than studying, maximizing enrichment requires helping the user avoid the platform altogether by recommending minimally tempting content or no recommendations at all. This illustrates how optimal recommendation strategies must dynamically adapt based on the relative enrichment value of users’ available alternatives, fundamentally challenging the one-size-fits-all approach of traditional recommender systems.
Our second contribution is to formulate a novel optimization objective for maximizing consumed enrichment and ensuring user well-being. In addition, we also prescribe the optimal recommendation strategy for this objective, with proof of optimality. Our model and optimization provide three key implications for design:
@@ 1.2. Optimizing Enrichment
    Consumption choices made by a user may prove to be a poor proxy for their end goal or true desire.
@@
    Items recommended do not necessarily equal items consumed, so it is not necessarily enough to recommend highly enriching items.
  3. (3)
    Recommending items can be costly. Tempting items may crowd out more enriching, but less tempting alternatives.
  4. (4)
    Maximizing enrichment requires understanding how enriching the outside option is, and whether the platform should be trying to retain the user while maximizing enrichment.
We note here that our model captures the settings where a user must choose which option to consume. Therefore, our results apply more to content platforms like YouTube and Netflix, than to social network platforms like X, Bluesky, Instagram, or Facebook, where the consumption of each item is nearly instantaneous and recommendation almost certainly implies consumption.
1.3. Learning from Data
Our proposed recommender algorithm requires knowledge of enrichment and temptation values for items. In practice, these are unobservable quantities that must be learned from data.
Existing literature offers several approaches to this challenge. Agarwal et al.(Agarwal et al., 2024) examine user return probability as a behavioral proxy for long-term satisfaction. Chang et al.(Chang et al., 2023) employ a latent variable model where past behaviors and context generate probability distributions over hidden user intents, which then predict future behavior. However, most prior studies combine solicited non-engagement signals with engagement metrics to infer user satisfaction. For instance, Cunningham et al.(Cunningham et al., 2024) propose using engagement diversity and item-level surveys. Milli et al.(Milli et al., 2021) demonstrate how to use specific observed behaviors as anchor variables to derive “value” rather than engagement, such as leveraging a “show me less of this” button. Other researchers suggest that using multiple behavioral signals rather than a single metric could provide better insights (Kleinberg et al., 2024). Indeed, Milli et al. (Milli et al., 2023b) show that up-weighting behaviors more indicative of whether users value specific content leads to improved “value” estimation in their model.
Our third contribution is a basic estimation framework that identifies enrichment and temptation from a combination of behavioral data and explicit user feedback about both on-platform items and outside options. Crucially, our framework imposes minimal, reasonable assumptions about users’ outside options, which are necessary for implementing the optimal recommender. This serves to demonstrate the feasibility of our proposed recommender in practical scenarios.
Our estimator is designed for our specific setting; however, in practice, ideas and insights from the aforementioned works could likely be combined with our model.
2. Behavioral Model
@@
We model user behavior within the confines of a single platform, considering $m$ users and $n$ items, where $m\gg n$. As previously discussed, for each user $j=1,\ldots,m$, each item $i=1,\ldots,n$ has two attributes: i) Enrichment $u_{j}(i)$ measures the inherent value the item provides to the user, and ii) Temptation $v_{j}(i)$ measures the appeal of the item to the user.
@@
where $f_{rating}$ is monotone and non-decreasing. Notice that our conception of rating data need not directly reveal a user’s enrichment and may also be influenced by the temptation of items.
An important assumption in our model is that $\lambda^{F}_{j}>\lambda^{C}_{j}\hskip 5.69054pt\forall j=1,\ldots,m$. This means that users tend to place more weight on enrichment when providing feedback than when choosing what to consume. This assumption is grounded in the observation that rating an item typically involves more deliberate reflection from a user compared to the often more impulsive act of choosing what to consume. This aligns with psychological theories suggesting that people are more likely to consider the inherent value when engaged in reflective thought processes.
3. Recommender Design
@@
We then induct on the number of rounds (the depth of the tree), $T$. The intuition for the proof here is that if a recommendation strategy $\mathcal{A}$ is optimal, it must be locally greedy after the first round because of the induction hypothesis. Then, we can either swap the item recommended by $\mathcal{A}$ in round 1 with a later item without losing enrichment, or we can replace the item recommended by $\mathcal{A}$ in round 1 with an item that provides better enrichment in expectation. A detailed proof is provided as auxiliary material. ∎
@@ 4. Model Estimation
The optimal recommender strategy proposed in section 3 requires knowledge of both on-platform items and off-platform outside options for each user, specifically their enrichment and temptation values, as well as user-specific parameters. However, in reality, platforms do not possess this comprehensive information and must instead estimate these components from available historical data. To address this challenge, we first extend our model to elucidate the determination of enrichment and temptation values for items. Subsequently, we provide a framework for estimating these critical components and generating recommendations that align with our proposed strategy. This estimation process bridges the gap between theoretical optimality and practical implementation, demonstrating that platforms can in fact leverage our recommendation strategy effectively, even in the absence of perfect information. By doing so, we not only enhance the applicability of our model but also provide valuable insights into user behavior and content characteristics.